I have a little doubt about Interest Table

i just want to know that how to calculate the present/future value/interest rate without Using the table .eg :

Plz hep me and EXPLAIN WITH THE BEST OF YOUR ABILITY .Plz solve the Example without using interest Table which will help me to determine/ find the initial Solution for my Doubt with a help!

For eg:-Please explain to the best of your ability
Answer #1

Solution 3-26:

Given that PV = $15,000, I = 4%/4 = 1% and n = 5*4 = 20

PV = A (PVIFA @ I, n)

$15,000 = A (PVIFA @ 1%, 20)

$15,000 = A [(1.01^20-1)/(0.01*1.01^20)]

$15,000 = A (18.04555)

A = $15,000/18.04555

A = $831 per quarter

Solution 3-27:                 

Given that A = $800, I = 5%/4 = 1.25% and n = 7*4 = 28

FV = A (FVIFA @ I, n)

FV = $800 (FVIFA @ 1.25%, 28)

FV = $800 [(1.0125^28-1)/0.0125]

FV = $800 (33.27938)

FV = $25,824

Solution 3-28:

a. PV =A (PVIFA @ I, n)

$1500 = A (PVIFA @ 18%/12, 3*12)

$1500 = A (PVIFA @ 1.5%, 36)

$1500 = A [(1.0015^36-1)/(0.0015*1.0015^36)]

$1500 = A (27.66068)

A = $54.30 per month

b. Interest payment = Principal x Interest rate

Interest payment = 1,500 x 0.015

Interest payment = $22.50

c. PV =A (PVIFA @ I, n)

PV = $54.30 (PVIFA 1.5%, 16)

PV = $54.30 [(1.015^16-1)/(0.015*1.015^16)]

PV = $54.30 (14.13126)

PV = $767.31